# Safe Interaction via Monte Carlo Linear-Quadratic Games

### v1

@@ V Simulations
In theory, our proposed MCLQ method converges towards robust, game-theoretic behaviors while offering real-time performance. Here we test both of these theoretical claims by comparing our approach to exact Hamilton-Jacobi methods and tractable LQ approximations. More specifically, we perform controlled experiments where simulated agents interact in three environments: point-mass, driving, and manipulator (see Figure 2). Each environment contains two agents. The simulated robot is attempting to complete its task (e.g., reaching a goal) while simulated humans move in its proximity. The robot must avoid collisions with these simulated humans — even when the humans take unexpected or noisy actions. A repository of our code is available here.
Independent Variables. We vary the robot’s controller across four levels. To test LQ approximations, we include NPG [15] and ILQ [4] approaches. These methods iteratively approximate the dynamics and cost as a LQ system and solve for the resulting Nash Equilibrium. To test exact solutions, we next solve for the true Nash Equilibrium (NE) using the Bellman Equation in Equation (7). Finally, we evaluate our MCLQ approach from Algorithm 1.
Environments. Below we describe the three simulated environments. All simulations were performed on an AMD Ryzen 7000 Series 5 CPU.
Point-Mass. Here the simulated human and robot move in a $2$D plane. The initial positions and velocities of both agents are randomized; across an interaction of $T=30$ timesteps, the robot attempts to reach a fixed goal location. Both agents have linear dynamics such that $x^{t+1}=x^{t}+u^{t}+w^{t}$, and the robot’s cost is quadratic (considering both distance to goal and distance to human). Because this environment is a linear-quadratic game, we anticipate that LQ approximations should find the Nash Equilibrium.
Driving. Our driving environment is taken from the iLQ baseline [4]. As before, one simulated human and robot move in a $2$D plane with randomized initial configurations. The dynamics of each agent follow the nonlinear bicycle model from [23]. Here the robot’s cost function is not quadratic:
@@
Manipulator. Our final environment contains two $7$-DoF robot arms in PyBullet. We treat one of these arms as the simulated human, and the other as the robot. Both agents are trying to pick up and place objects within a shared workspace. The dynamics of the arms and the task are nonlinear, and the system state $x\in\mathcal{X}$ is $26$-dimensional (including manipulators and objects). We leverage a cost function similar to driving. The robot moves to reach, grasp, and place items while avoiding collisions with the other arm.
@@
$\begin{split}\pi_{\mathcal{R}}\left(w^{t}\mid x^{t}\right)&\propto\exp\left(-\alpha\cdot J_{H}(x^{t},w_{i}^{t})\right)\\ \text{s.t.}\leavevmode\nobreak\ x^{t+1}&=f(x^{t},u^{t_{0}},w^{t})\end{split}$
Increasing $\alpha\rightarrow\infty$ causes the human to always take the optimal action, while decreasing $\alpha\rightarrow 0$ causes the human to act randomly [7]. In our simulations we set $\alpha=7.5$. When computing the cost for future timesteps the simulated human assumes the robot will repeat its most recent action $u^{t_{0}}$ (e.g., the robot will keep moving with the same velocity). In each environment the human had a task that was independent of the robot’s objective — for instance, in driving the human tried to reach their own goal position.
Results. The results from our first simulation are summarized in Figure 2. Across all three environments, MCLQ achieved costs that were closest to the Nash Equilibrium (NE) while also requiring the least amount of computation time. In point-mass we highlight that all methods reached similar cost; this matches our expectations, because point-mass was an LQ system. By contrast, in non-LQ systems (driving and manipulator) the LQ approximations made by NPG and ILQ fell short, leading to suboptimal performance. Computing the exact solution with NE was time-consuming and not always feasible: indeed, in the $26$-dimension manipulator task, we were unable to compute the true NE because of the high-dimensional and continuous state-action space.
Figure 3: Simulation results for a modified point-mass environment where we adjust the safety margin $\lambda$ in MCLQ. Increasing $\lambda$ causes the MCLQ robot to consider a wider range of worst case human actions, resulting in more conservative behavior. Conversely, decreasing $\lambda$ causes the MCLQ robot to increasingly rely on its nominal human model. Unlike LQ approximations, our proposed method gives designers the flexibility to tune $\lambda$ and adjust the safety margin.
Adjusting Safety Margins. As discussed in Section IV-D, one feature of our approach is the designer-selected safety margin. MCLQ robots reason over the worst-case human action within bounds $\lambda$. By increasing $\lambda$ in Equation (18) the designer makes the robot more risk-averse (i.e., the robot considers larger deviations from the nominal human model). Conversely, decreasing $\lambda$ makes the robot more risk-seeking (i.e., the robot increasingly relies on its human model). Within our simulations from Figure 2 we left this value of $\lambda$ fixed at risk-neutral behavior. Now we explore how increasing and decreasing $\lambda$ changes the robot’s performance in a modified point-mass environment.
@@ VI User Study
[…]
Our simulations from Section V support our theoretical analysis, and suggest that MCLQ converges towards maximally-robust behaviors while minimizing computation time. We next evaluate our method in a real-world setting with $N=24$ in-person human participants. Here users walk around a room to complete an assembly task while an autonomous drone circles that room to inspect the parts (see Figure 1). The robot modifies its high-level trajectory to avoid getting to close to the human workers. We compare two real-time safety filters for adjusting the drone’s behavior: ILQ [4] and MCLQ. We selected ILQ here because it was the best performing real-time baseline from our simulations. Videos of our user study are available here.
Experimental Setup. Participants interacted with a Crazyflie 2.1+ (Bitcraze) during the assembly task. The participant’s objective was to construct a Lego tower at the central station. The blocks needed for building that tower were scattered in three other stations located along the perimeter of the workspace. Accordingly, users needed to move back and forth through the workspace to acquire blocks and build their tower. The drone’s objective was to monitor the workspace during the task by completing as many revolutions around the central station as possible. As the drone moved around the central station it repeatedly intersected with the human worker: at these intersections the drone should to take actions to help avoid the moving participant.
Participants and Procedure. We recruited $24$ participants ($6$ female, average age $24.5\pm 4.5$) for this user study. Of the $24$ participants, $8$ had not previously used drones and $5$ did not have prior experience with robotics. Participants received monetary compensation for their time and provided written consent according to university guidelines (IRB #23-1237).
@@
Dependent Measures — Objective. We using tracking devices (VIVE) to measure the states and actions of the drone and human at each timestep of the interaction. To assess objective performance, we considered two metrics. For our first metric (collisions) we counted the number of times the drone was within a radius of $0.5$m from the human. Our second metric (revolutions) is the number of revolutions the drone completed within the five minute trial. Lower values for collisions indicate that the drone is maintaining safety, while higher values for revolutions show that the drone is performing the task more efficiently.
@@ VII Conclusion
In this paper we presented a real-time safety filter for human-robot interaction. Our approach ensures that the robot is robust to noisy and unexpected human behaviors within designer-specified bounds. To achieve this game-theoretic safety with tractable computation, we combined linear-quadratic approximations with stochastic local searches. Our theoretical analysis showed the resulting MCLQ algorithm converges towards the ideal Nash Equilibrium policy while providing flexible and efficient implementation. Across multiple simulations and a user study, we observed that MCLQ advances both objective and subjective safety measures when compared to state-of-the-art control alternatives.

### v2

@@ V Simulations
In theory, our proposed MCLQ method moves towards robust, game-theoretic behaviors while offering real-time performance. Here we test both of these theoretical claims by comparing our approach to exact Hamilton-Jacobi methods and tractable LQ approximations. More specifically, we perform controlled experiments where simulated agents interact in three environments: point-mass, driving, and manipulator (see Figure 2). Each environment contains two agents. The simulated robot is attempting to complete its task (e.g., reaching a goal) while simulated humans move in its proximity. The robot must avoid collisions with these simulated humans — even when the humans take unexpected or noisy actions.
Independent Variables. We vary the robot’s controller across four levels. To test LQ approximations, we include DARE-nA [wu2023learning] and ILQGames [fridovich2020efficient] approaches. These methods iteratively approximate the dynamics and cost as a LQ system and solve for the resulting Nash Equilibrium. We used the official repositories for DARE-nA and ILQGames when applicable. To test exact solutions, we next solve for the true Nash Equilibrium (NE) using the Bellman Equation. Finally, we evaluate our MCLQ approach from Section IV.
Environments. Below we describe the three simulated environments. All simulations were performed on an AMD Ryzen 7000 Series 5 CPU with multithreading enabled.
Point-Mass. Here the simulated human and robot move in a 22D plane. The initial positions and velocities of both agents are randomized; across an interaction of $T=30$ timesteps, the robot attempts to reach a fixed goal location. Both agents have linear dynamics such that $x^{t+1}=x^{t}+u^{t}+w^{t}$, and the robot’s cost is quadratic (considering both distance to goal and distance to human). Because this environment is a linear-quadratic game, we anticipate that LQ approximations should find the Nash Equilibrium.
Driving. Our driving environment is taken from the iLQ baseline [fridovich2020efficient]. As before, one simulated human and robot move in a 22D plane with randomized initial configurations. The dynamics of each agent follow a nonlinear bicycle model. Here the robot’s cost function is not quadratic:
@@
Manipulator. Our final environment contains two $7$-DoF robot arms in PyBullet. We treat one of these arms as the simulated human, and the other as the robot. Both agents are trying to pick up and place objects within a shared workspace. The dynamics of the arms and the task are nonlinear, and the system state $x\in\mathcal{X}$ is $26$-dimensional (including manipulators and objects). We leverage a cost function similar to driving. The robot moves to reach, grasp, and place items while avoiding collisions with the other arm. We separate this task into two separate sub-tasks: reaching and placing, each with separate cost functions. Settings with grasping violate the smooth dynamics assumptions of Section III. To still use LQ approximations for the environment, the sub-task is automatically switched when the reaching sub-task’s cost is below a certain threshold.
@@
$\begin{split}\pi_{\mathcal{R}}\left(w^{t}\mid x^{t}\right)&\propto\exp\left(-\alpha\cdot J_{H}(x^{t},w_{i}^{t})\right)\\ \text{s.t.}\penalty 10000\ x^{t+1}&=f(x^{t},u^{t_{0}},w^{t})\end{split}$
Increasing $\alpha\rightarrow\infty$ causes the human to always take the optimal action, while decreasing $\alpha\rightarrow 0$ causes the human to act randomly [fridovich2020confidence]. In our simulations we set $\alpha=7.5$. When computing the cost for future timesteps the simulated human assumes the robot will repeat its most recent action $u^{t_{0}}$ (e.g., the robot will keep moving with the same velocity). In each environment the human had a task that was independent of the robot’s objective — for instance, in driving the human tried to reach their own goal position.
Results. The results from our first simulation are summarized in Figure 2. Across all three environments, MCLQ achieved costs that were closest to the Nash Equilibrium (NE) while also requiring the least amount of computation time. In point-mass we highlight that all methods reached similar cost; this matches our expectations, because point-mass was an LQ system. By contrast, in non-LQ systems (driving and manipulator) the LQ approximations made by DARE-nA and ILQ fell short, leading to suboptimal performance. Computing the exact solution with NE was time-consuming and not always feasible: indeed, in the $26$-dimension manipulator task, we were unable to compute the true NE because of the high-dimensional and continuous state-action space.
Figure 3: Simulation results for a modified point-mass environment where we adjust the safety margin $\lambda$ in MCLQ. Increasing $\lambda$ causes the MCLQ robot to consider a wider range of worst case human actions, resulting in more conservative behavior. Conversely, decreasing $\lambda$ causes the MCLQ robot to increasingly rely on its nominal human model. Unlike LQ approximations, our proposed method gives designers the flexibility to tune $\lambda$ and adjust the safety margin. In practice, this results in risk-averse robots that avoid all possible collisions but reach the goal more slowly, or risk-seeking robots that allow some potential collisions and rapidly move to the goal.
Adjusting Safety Margins. As discussed in Section IV-C, one feature of our approach is the designer-selected safety margin. MCLQ robots reason over the worst-case human action within bounds $\lambda$. By increasing $\lambda$ in Equation (19) the designer makes the robot more risk-averse (i.e., the robot considers larger deviations from the nominal human model). Conversely, decreasing $\lambda$ makes the robot more risk-seeking (i.e., the robot increasingly relies on its human model). Within our simulations from Figure 2 we left this value of $\lambda$ fixed at risk-neutral behavior. Now we explore how increasing and decreasing $\lambda$ changes the robot’s performance in a modified point-mass environment.
@@ VI User Study
Figure 4: Results from our user study in Section VI. Participants walked around a room to assemble a tower; a drone completed revolutions around the same workspace to monitor the human’s progress (also see Figure 1). (Left) The average number of revolutions the drone completed and the average number of collisions. Here “collisions” occurred when the drone was within $0.5$ meters of the human. The proposed MCLQ algorithm adjusts the robot’s behavior to increase safety (fewer collisions) while also enhancing performance (more revolutions). (Right) After interacting with each algorithm participants answered survey questions about how safe, predictable, and attentive the robot was. Ratings suggest that participants perceived MCLQ to be a safer system. Error bars show standard deviation and an * denotes statistical significance ($p<.05$).
Our simulations from Section V support our theoretical analysis, and suggest that MCLQ converges towards maximally-robust behaviors while minimizing computation time. We next evaluate our method in a real-world setting with $N=24$ in-person human participants. Here users walk around a room to complete an assembly task while an autonomous drone circles that room to inspect the parts (see Figure 1). The robot modifies its high-level trajectory to avoid getting to close to the human workers. We compare two real-time methods for safe interactions that adjust the drone’s behavior: ILQ [fridovich2020efficient] and MCLQ. We selected ILQ here because it was the best performing real-time baseline from our simulations.
Experimental Setup. Participants interacted with a Crazyflie 2.1+ (Bitcraze) during the assembly task. The participant’s objective was to construct a Lego tower at the central station. The blocks needed for building that tower were scattered in three other stations located along the perimeter of the workspace. Accordingly, users needed to move back and forth through the workspace to acquire blocks and build their tower. The drone’s objective was to monitor the workspace during the task by completing as many revolutions around the central station as possible. As the drone moved around the central station it repeatedly intersected with the human worker: here the drone should take actions that avoid the moving participant. Participants interacted with the drone separately (i.e., the drone interacted with one user at a time).
Participants and Procedure. We recruited $24$ participants ($6$ female, age $24.5\pm 4.5$) for this user study. Of the $24$ participants, $8$ had not used drones and $5$ did not have experience with robotics. Participants received monetary compensation for their time and provided written consent according to university guidelines (IRB #23-1237).
@@
Dependent Measures — Objective. We using tracking devices (VIVE) to measure the states and actions of the drone and human at each timestep of the interaction. To assess objective performance, we considered two metrics. For our first metric (collisions) we counted the number of times the drone was within a radius of 0.50.5m from the human. Our second metric (revolutions) is the number of revolutions the drone completed within the five minute trial. Lower values for collisions indicate that the drone is maintaining safety, while higher values for revolutions show that the drone is performing the task more efficiently.
@@ VII Conclusion
In this paper we presented a real-time safety approach for human-robot interaction. Our approach ensures that the robot is robust to noisy and unexpected human behaviors within designer-specified bounds. To achieve this game-theoretic safety with tractable computation, we combined linear-quadratic approximations with stochastic local searches. Our theoretical and empirical analysis showed the resulting MCLQ algorithm converges towards robust robot policies while providing flexible and efficient implementation. Across multiple simulations and a user study, we observed that MCLQ advances both objective and subjective safety measures when compared to state-of-the-art control alternatives.